An urn contains 7 blue, 7 yellow, and 8 gray marbles. Two marbles are randomly drawn from the urn.
Find the probability the two marbles are the same color.

Found 2 solutions by VFBundy, Boreal:
Answer by VFBundy(438)
Answer by Boreal(15235)

22 marbles and choose 2.
There are 22C2 or 231 ways.
for 2 blue or 2 yellow, it is (7/22)(6/21)=42/462 each or 42/231
for 2 gray it is (8/22)(7/21)=56/432 or 28/231
So the answer is 70/231 or 0.3030